Can you please help me find the intercepts of a line that contains the coordinate points of (-7, 12) and (14, -6) ? I know how to get the slope, which would be -6/7 , but where do I go from there? Thank you !
Answer by richwmiller(17219)

y-y1=m*(x-x1)
put in either point
y-12=-6/7*(x-(-7))
y = -(6 x)/7+6
y = 6-(6x)/7